Match the graphs with the functions. a 10 a - 10 10 fr ns -n Te 10 -10 -10 Drag your answer(s) to the correct position
Answer #1

0 y = |x+2. (2) 9 е – 6х2 ry X (0,22 . 7x. Гу: - 6x2 9=-2 |x| T + 8 = -(х+4)- 3 9 (-6) 7x 9 = 2 )) ,-36) S 9 = 1х -2 9= -(х+4५ = (x-5. (5) Y ५-12+2] (6,30) (-2,0) न (6,0) C2,00 (10) ५ ५-2121 m Y-6x2 + x 가 (2- 2+--- -8 (هر 6) (-6,०)